OpenStudy (anonymous):

Absolute max/min! f(x)=1-7x^2, on the interval -3 less than or equal to x less than or equal to 2.

OpenStudy (anonymous):

max is at x = 0 (so max value is 1) check the endpoints for the min

OpenStudy (anonymous):

that is, check \[f(-3)\] and \[f(2)\] whatever is smallest is smallest. i will do doubt be \[f(-3)\] since -3 is further from 0 than 2 is

OpenStudy (anonymous):

well I had the max no problem, but I can't figure out the min value

OpenStudy (anonymous):

those you have to check

OpenStudy (anonymous):

Yeah, I can't seem to solve it. Thank you though

OpenStudy (anonymous):

\[f(-3)=1-7\times 9=-62\] is the min

OpenStudy (anonymous):

because \[f(2)=1-7\times 4=-27\] and \[-62<-27\] so \[-62\] is the mininum value on that interval

OpenStudy (anonymous):

Thank you so much! :D you are a life saver